2. (5) There is a sub-saharan African community with 10,000 individuals. If 9% of this population is born with a severe form of sickle-cell anemia (ss), determine the genotype and allelic frequencies in the population and tell me how many individuals would be expected to be SS, Ss and ss (Im just asking you to fill in the tables): Allelic Frequencies: Frequency of S Frequency of s Genotypic Frequencies: Frequency of SS Frequency of Ss Frequency of ss How many people? How many people? How many people?
Answer #1

Allele frequency estimation:
Step 1: Recesseive allele frequnecy estimation:

Sickle cell anemia is caused by homozogygous recesseive condition.

The precentage of sickle cell population = 9%, the frequency = 9X10000/100 = 900.
so, ss or s^2= 900, the proportion of ss pr s^2 = 900/10000 = 0.09, so s= sqrt(0.09) = 0.3.

The frequency of s = 0.3.

Step:2: Dominant allele frequency estimaton:
As the frequency of S + s = 1; S= 1-s => 1-0.3 = 0.7.

The frequency of S = 0.7.


Geneotypic frequenceis:
SS = 0.7X0.7 = 0.49. The people = 0.49X10000 = 4900
Ss = 2x0.7 x 0.3 = 0.42 The people = 0.42x10000 = 4200
ss = 0.3x0.3 = 0.09. The people = 0.09x10000 = 900